Write the slope-intercept form of the equation of the lne shown in the graph. Please show work. (5,1) (-5,-5)


Can someone please help me, I just dont understand this at all. Any help I can get will be greatly appreciated..Thank you so much for your time and efforts.

Answer by oscargut(2103)

Write the slope-intercept form of the equation of the lne shown in the graph. Please show work. (5,1) (-5,-5)
Solution:
Slope = (-5-1)/(-5-5) = -6/-10 = 3/5
So equation is: y =(3/5)x+b (we have to find b now)
Since (5,1) is in the line:
1=(3/5)(5)+b
1 = 3+b
b = -2
Answer: y = (3/5) x -2
If you have any doubt or if you have more problems my e-mail is:
mthman@gmail.com